### v1

# FILP-3D: Enhancing 3D Few-shot Class-incremental Learning with Pre-trained Vision-Language Models

Abstract
Few-shot class-incremental learning (FSCIL) aims to mitigate the catastrophic forgetting issue when a model is incrementally trained on limited data. While the Contrastive Vision-Language Pre-Training (CLIP) model has been effective in addressing 2D few/zero-shot learning tasks, its direct application to 3D FSCIL faces limitations. These limitations arise from feature space misalignment and significant noise in real-world scanned 3D data. To address these challenges, we introduce two novel components: the Redundant Feature Eliminator (RFE) and the Spatial Noise Compensator (SNC). RFE aligns the feature spaces of input point clouds and their embeddings by performing a unique dimensionality reduction on the feature space of pre-trained models (PTMs), effectively eliminating redundant information without compromising semantic integrity. On the other hand, SNC is a graph-based 3D model designed to capture robust geometric information within point clouds, thereby augmenting the knowledge lost due to projection, particularly when processing real-world scanned data. Considering the imbalance in existing 3D datasets, we also propose new evaluation metrics that offer a more nuanced assessment of a 3D FSCIL model. Traditional accuracy metrics are proved to be biased; thus, our metrics focus on the model’s proficiency in learning new classes while maintaining the balance between old and new classes. Experimental results on both established 3D FSCIL benchmarks and our dataset demonstrate that our approach significantly outperforms existing state-of-the-art methods. Code is available at Code is available at https://github.com/HIT-leaderone/FLIP-3D
1 Introduction
The proliferation of 3D content in recent years has been marked by both synthetic creations [2, 5] and real-world reconstruction [29, 24, 35]. This expansion in data scale naturally introduces new classes into the 3D domain. However, these emerging categories often comprise a limited number of instances, posing a challenge for existing 3D recognition models. Consequently, 3D few-shot class-incremental learning (FSCIL) has become increasingly important in practical applications.
Previous work in 2D few-shot and zero-shot tasks [20, 47] has shown that Pre-Training Models (PTMs) excel in incremental learning scenarios, often outperforming non-PTM-based approaches. This superior performance is largely attributed to the prior knowledge that PTMs acquire [6, 10, 21], which enhances generalization in downstream tasks. Motivated by these successes, we aim to leverage PTMs to imbue our model with shape-related prior knowledge. However, the limited scale of available 3D data hampers the effectiveness of existing 3D PTMs in downstream tasks. To address this, recent studies [43, 11, 40] have successfully aligned 3D representations with Vision-Language (V-L) PTM knowledge. Building on this, we map point cloud data into a cross-modal space, utilizing vision and language pre-training to improve performance in 3D FSCIL tasks. Specifically, we generate multi-view depth maps from point clouds and employ template text prompts for classification, such as “an image of a {class name}”.
Figure 1: CLIP learns a large amount of prior knowledge from massive image-text pairs. Thus pre-aligned image and text features contain sufficient shape-related prior knowledge. Along with the elimination of redundant information (RFE) and the compensator of 3D fine-grained information (SNC), the performance in 3D FSCIL can be significantly improved.
While the integration of Vision-Language (V-L) Pre-Training Models (PTMs) into FSCIL tasks has yielded improvements over previous methods, two key challenges remain to limit performance. First, V-L PTMs are designed to capture detailed visual features such as color and texture. However, point cloud data, being a collection of discrete points, lacks the information to represent these intricate visual details. This discrepancy can lead to inaccurate feature extraction, adversely affecting performance. Second, the use of V-L PTMs amplifies the issue of noise sensitivity, particularly in real-world scanned data. Consequently, classification performance suffers significantly if noise disrupts the rendering of depth maps and object contours.
To tackle the aforementioned challenges, we introduce Few-shot class Incremental Learning tasks with Pre-training on 3D (FILP-3D), a framework that employs CLIP as its backbone and incorporates two innovative components: Redundant Feature Eliminator (RFE) and Spatial Noise Compensator (SNC), as is shown in Figure 1. RFE serves as a specialized dimensionality reduction technique, designed to remove redundant features while preserving semantic content. By precisely compressing dimensions within the V-L PTMs, RFE facilitates better alignment between the feature spaces of point clouds and their corresponding embeddings. On the other hand, SNC is a graph-based 3D model tailored to extract robust geometric information from point clouds. This additional layer of information enhances the geometric relationships between object components, mitigating the loss of knowledge that occurs during projection, especially in noisy, real-world scanned data. As a result, our FILP-3D framework gains improved resilience against noise interference.
Furthermore, we observe that the numbers of samples for different classes in test datasets can vary by up to two orders of magnitude, leading to a significant imbalance. This imbalance skews traditional accuracy metrics, making them less reliable for comprehensive evaluation. To address this, we introduce new evaluation metrics, namely $\rm NCAcc$ and $\rm F_{FSCIL}$. These metrics are designed to assess the model’s ability to learn new classes effectively while maintaining a balance between old and new classes, thereby providing a more nuanced evaluation of the model performance.
In a nutshell, our contributions are three-fold:
  • We pioneer the application of Vision-Language Pre-Training Models (V-L PTMs) to 3D Few-Shot Class-Incremental Learning (FSCIL) tasks, achieving performance gains over existing models. The general embedding provided by V-L PTM, along with its embedded prior knowledge, can complement the missing information in few-shot tasks and alleviate catastrophic forgetting during the incremental learning process.
  • We introduce FILP-3D, a framework that incorporates two innovative modules: the Redundant Feature Eliminator (RFE) and the Spatial Noise Compensator (SNC). RFE specifically addresses feature space misalignment, while SNC is designed to mitigate the adverse effects of noise on the model. FILP-3D yields notable performance improvements, especially in metrics associated with novel classes.
  • We develop new metrics, namely $\rm NCAcc$ and $\rm F_{FSCIL}$ to provide a more nuanced evaluation. These metrics assess a model’s ability to adeptly learn new classes without compromising the performance on existing classes, offering a more comprehensive evaluation framework for 3D FSCIL tasks.
2 Related Work
2.1 3D Point Cloud understanding
In recent years, many works have been proposed to classify 3D point cloud objects. PointNet [18] and PointNet++ [19] design the architecture to maintain natural invariances of the data. DGCNN [34] connects the point set into a graph and designs a local neighbor aggregation strategy.
Some works employ masked point modeling [7, 12] as a 3D self-supervised learning strategy to achieve great success. For example, Point BERT [39] uses a pre-trained tokenizer to predict discrete point labels, while Point MAE [17] and Point-M2AE [42] apply masked autoencoders to directly reconstruct the masked 3D coordinates.
[…]
2.2 Few-Shot Class-Incremental Learning
The FSCIL problem was proposed by Tao et al. [28]. Concretely, FSCIL aims at learning from severely insufficient samples incrementally while preserving already learned knowledge. TOPIC [28] uses a neural gas network to learn and preserve the topology of features. Subsequently, CEC [41] utilizes a graph model to propagate context information between classifiers for adaptation. Also, some models like FACT [46] try to use virtual prototypes to reserve for new ones, thus ensuring incremental learning ability. The SOTA method BiDist [45] utilizes a novel distillation structure to alleviate the effects of forgetting.
Recently, with the breakthroughs in 2D PTMs, a number of works attempt to leverage the vast knowledge acquired by 2D PTMs, which is effective in learning new concepts and alleviating the problem of forgetting [22, 47, 33]. With almost perfect performance, these models have attracted a lot of interest and attention. The preceding methods are all 2D methods, except for Chowdhury et al. [3]’s work which explores the FSCIL task in 3D point cloud data. However, Chowdhury et al. [3]’s work neither supplements knowledge for few-shot data nor addresses the high-noise nature of real-world scanned data, resulting in limited performance in 3D FSCIL tasks. In contrast, our FILP-3D successfully addresses the shortcomings mentioned above by introducing V-L PTMs and two newly proposed modules.
3 Proposed Method
3.1 Problem Formulation
Assuming a sequence of $B$ tasks $\mathcal{D}=\{\mathcal{D}^{1},\mathcal{D}^{2},\dots,\mathcal{D}^{B}\}$, FSCIL methods incrementally recognize novel classes with a small amount of training data. For the $b$-th task $\mathcal{D}^{b}=\{(\mathbf{x}_{i}^{b},y_{i}^{b})\}_{i=1}^{n_{b}}$, we have $n_{b}$ training samples, in which an instance $\mathbf{x}_{i}^{b}$ has a class label $y_{i}^{b}$ and $y_{i}^{b}$ belongs to the label set $Y_{b}$. We stipulate that $Y_{b}\cap Y_{b^{\prime}}=\emptyset$ when $b\not=b^{\prime}$. After the $b$-th training task, trained models are required to classify test sets of all the previous tasks $\{\mathcal{D}^{1},\dots,\mathcal{D}^{b}\}$. Note that, in the 3D FSCIL setting, $\mathbf{x}_{i}^{b}\in\mathbb{R}^{P\times 3}$ denotes a point cloud object, where $P$ donates the number of points in a point cloud object. $\mathcal{D}^{1}$ is the base task with a large-scale 3D training dataset, while much fewer samples are included in the incremental task $\mathcal{D}^{b}$, i.e., $n_{b}\ll n_{1}$ for $b>1$.
3.2 Model Overview
Figure 2: Overview of FILP-3D. FILP-3D mainly consists of three components, i.e., 3D branch (SNC), 2D branch, and textual prototypes. Then, the 3D feature and the 2D feature will be fused as a global feature, and be used to calculate probability through the REF module alongside textual prototypes.
The framework of FILP-3D is shown in Figure 2. For each training sample $(\mathbf{x}_{i}^{b},y_{i}^{b})$, we mainly have three branches: 1) point cloud $\mathbf{x}_{i}^{b}$ is fed into 3D encoder to obtain an original point feature $\mathbf{f}^{3D}$ and then aligned to $\mathbf{f}^{p}$, 2) $\mathbf{x}_{i}^{b}$ is then rendered as multi-view depth maps ${\rm\mathbf{D}}^{b}_{i}$, embedded as depth features ${\rm\mathbf{F}}^{2D}$, and finally merged into a global depth feature $\mathbf{f}^{d}$, 3) visible class names are templated into text prompts and encoded as text features ${\rm\mathbf{F}}^{t}$. The point features $\mathbf{f}^{p}$ and the global depth feature $\mathbf{f}^{d}$ are then fused as a global feature $\mathbf{f}^{g}$. Afterward, we use pre-processed principal components $\mathbf{V}$ (refer to Sec. 5 in Suppl. for more details of pre-processing) to eliminate redundant dimensions of $\mathbf{f}^{g}$ and $\mathbf{F}^{t}$ and generate $\widetilde{\mathbf{f}}^{g}$ and $\widetilde{\mathbf{F}}^{t}$. The final predicted probability is calculated by our proposed renormalized cosine similarity.
3.3 PTM is a Good 3D FSCIL Learner
Recent works [20, 47] have shown that PTMs substantially enhance the performance in incremental learning tasks. However, Chowdhury et al. [3]’s method, in the absence of shape-related prior knowledge, struggles with the challenge of catastrophic forgetting in 3D continual learning. Further complicating matters, the point features extracted by PointNet [18] are not in alignment with the prototypes generated through word2vec [15]. Thus, Chowdhury et al. [3]’s work finds it challenging to simultaneously retain features associated with base classes (representing old knowledge) and integrate features of novel classes (representing new knowledge).
In response to the aforementioned challenges, our initial approach was to embed shape-related prior knowledge via a 3D PTM. However, the substantial disparity in quality and volume between 3D data and its text/2D counterparts renders current 3D PTMs less effective in generalizing to downstream tasks. Motivated by this limitation, we turn to an alternative method. Drawing inspiration from recent studies [43, 11], we leverage CLIP [21] to indirectly infuse shape-related prior knowledge through projection. We christen this new framework as SimpleCIL-3D.
Specifically, we project the point cloud data into multi-view depth maps $\mathbf{D}^{b}_{1:N}$. A pre-trained ViT [8] in CLIP is then deployed to extract depth features $\mathbf{F}^{2D}\in\mathbb{R}^{N\times C}$. Here, $N$ is the number of views, and $C$ is the embedding dimension of ViT. To allow incremental tasks, a learnable merger is attached, formulating global depth features $\mathbf{f}^{d}\in\mathbb{R}^{C}$ as follows,
$\mathbf{f}^{d}=f^{1}_{d}({\rm ReLU}(f^{2}_{d}({\rm concat}(\mathbf{F}_{1:N}^{2D}))))$ (1)
where $f^{1}_{d}$, $f^{2}_{d}$ are two learnable MLPs.
The efficacy of setting classifier weights to average embeddings (referred to as prototypes) for CIL tasks has been well-established by [25]. However, while Chowdhury et al. [3]’s method employs word2vec to construct prototypes for new classes, it falls short in encapsulating the average semantics of 3D point clouds. In our approach, the depth features, denoted as $\mathbf{f}^{d}$, are pre-aligned with text embeddings of CLIP. This enables SimpleCIL-3D to progressively produce prototypes using a CLIP text encoder. For each class labeled as $k$ with its respective name $t_{k}$, we introduce a template text prompt: “an image or projection or sketch of a $t_{k}$”. This is mapped to the CLIP prototype symbolized by $\mathbf{F}^{t}_{k}\in\mathbb{R}^{C}$. Owing to the inherent association of our method between image and textual representations, there is no compulsion to realign the two modalities during incremental phases, unlike strategies such as in Chowdhury et al. [3]’s work. As a result, we can directly utilize the cosine similarity between $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$ to represent the logit for class $k$. The ultimate probability prediction, represented by $\mathbf{p}$, is formulated as follows:
$l_{k}=\cos(\mathbf{f}^{d},\mathbf{F}^{t}_{k}),\quad\mathbf{p}={\rm softmax}([l_{1},\dots,l_{K}]).$ (2)
where $\cos(\cdot,\cdot)$ denotes the cosine similarity, and $K$ is the number of visual classes.
3.4 Ensure Performance of PTM in 3D FSCIL
Large-scale V-L PTMs undeniably serve as rich repositories of prior knowledge. However, their application to 3D FSCIL tasks reveals unique challenges. First, point cloud data, which consists of discrete points in 3D space, primarily represents geometric features of objects. It inherently misses richer visual nuances, such as color and texture. This deficiency is especially significant given CLIP’s visual encoder, which inherently seeks and incorporates these visual details into its semantic representation. Consequently, this could jeopardize the precise classification of depth maps. Second, point clouds derived from real-world scans often suffer from noise, which distorts the extraction of depth-centric features. This problem is accentuated in multi-view projection techniques, where a significant fraction of the points might be occluded, thereby increasing the potential noise. Such disturbances can upset the balance in few-shot incremental learning, leading to overfitting. Addressing these challenges is imperative when tailoring PTMs for 3D FSCIL tasks. In the ensuing sections, we aim to enhance SimpleCIL-3D with elucidate solutions tailored to these specific hurdles, to obtain FILP-3D.
3.4.1 Redundant Feature Eliminator
We propose our design to mitigate the adverse effect of superfluous features here. Our empirical observations indicate that redundant features typically exhibit minimal inter-class distinction, leading to constrained variance. By contrast, semantic features pertinent for classification distinctly express a pronounced bias for each class, resulting in a more pronounced variance.  Chowdhury et al. [3], Zhu et al. [48] also yield empirical observations akin to ours. Drawing upon this insight, we can discern between redundant and semantic features based on their variance. Subsequently, we employ Dimensionality Reduction (DR) techniques to preserve the semantic essence while condensing the extraneous features.
Figure 3: The current feature space is generated by three principal components. Each dimension’s one-hot vector represents a principal component. The first two dimensions contain semantic information (high variance in the x and y axes), while the third dimension serves as a redundant component (low variance in the z-axis). By transforming feature vector (green) into the feature space mentioned above, we can notice that projection can eliminate redundant information, while normalization will improperly stretch the semantic information.
Following the above discussions, we can obtain the following formulation for $\widetilde{l}_{k}$, i.e., the logit after eliminating superfluous features. For the redundant feature eliminated logit of the $k$-th class (the detail discussion is in Sec. 1 of Suppl.), we have:
$\widetilde{l}_{k}=\frac{\widetilde{\mathbf{f}}^{d}\widetilde{\mathbf{F}}^{t^{\mathrm{T}}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$ (3)
where $\widetilde{\mathbf{f}}^{d}=\mathbf{V}^{\mathrm{T}}\mathbf{f}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}=\mathbf{V}^{\mathrm{T}}\mathbf{F}^{t}_{k}$. Here, $\mathbf{V}=[\mathbf{v}_{i}]_{M}$ consists of the principal components extracted from the base task ($M$ is the number of principal components)
Note that, different from calculating the cosine similarity of $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$, the denominator in Eq. 3 is the norm of the original features $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$. Principal features are directly dot multiplied after normalizing with the original features, and such a procedure is named as the renormalized cosine similarity (RCS). We can qualitatively compare the above two similarity formulations: If the mode lengths of $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$ are relatively small, it indicates that only a small portion of the features can be accurately extracted. Performing normalization under such circumstances will magnify the originally small features, resulting in incorrect classification, as shown in Figure 3. Conversely, the dot method has no such improper stretching operations.
3.4.2 Spatial Noise Compensator
SimpleCIL-3D unifies pre-aligned image and text encoders through comprehensive data training. This ensures a substantial reservoir of prior knowledge, minimizing the potential for information loss during any subsequent alignment processes. Given these attributes, it stands as a markedly superior choice when juxtaposed against partially-aligned 3D pre-trained models. Nonetheless, the projection-based methodology inherent to SimpleCIL-3D exhibits heightened sensitivity to both noise and viewpoint selection. When noise obscures the clarity of depth maps, rendering them ineffective in delineating object contours, there is a marked downturn in classification efficacy. As a countermeasure, we propose the incorporation of a graph-based 3D model. This addition is devised to augment the compromised information, bolstering our model’s resilience against noise disturbances.
The 3D module yields features that are invariant to transformations, thereby enhancing their robustness against noise. On the other hand, the multi-view model, fortified by PTM, extracts semantically richer and more encompassing information. To harness the strengths of both, we amalgamate the features derived from the 3D module and the multi-view model. The global feature is computed as:
$\mathbf{f}^{p}=f^{2}_{p}({\rm ReLU}(f^{1}_{p}(\mathbf{f}^{3D})))$ (4)
$\mathbf{f}^{g}=\frac{1}{2}({\rm max}(\mathbf{f}^{d},\mathbf{f}^{p})+{\rm avg}(\mathbf{f}^{d},\mathbf{f}^{p}))$ (5)
where $\mathbf{f}^{3D}$ is the feature extracted by 3D encoder from raw point cloud data $\mathbf{x}_{i}^{b}$, and $\mathbf{f}^{g}$ represents new global features used to replace $\mathbf{f}^{d}$ for further tasks. $f^{1}_{p}$, $f^{2}_{p}$ are learnable MLPs, so that 3D features $\mathbf{f}^{p}$ can be aligned to multi-view features $\mathbf{f}^{d}$.
Using this approach, any information inadequately captured due to viewpoint selection is supplemented by the 3D channel using the ${\rm max}$ operation. Concurrently, overly dominant features are moderated through the ${\rm average}$ operation for balance.
3.5 Training Objective
The parameters of the 3D encoder and the multi-view render & encoder are fixed during both base and incremental training stage. Only Merger (Eq. 1) for depth feature fusion and Adapter (Eq. 4) for 3D feature alignment are trained. The classification loss $\mathcal{L}^{b}_{cls}$ for $b$-th task can be calculated as follows:
$\mathcal{L}^{b}_{cls}=\frac{1}{|\mathcal{D}^{b}|}\sum_{i=1}^{|\mathcal{D}^{b}|}\mathcal{L}_{ce}(\mathbf{p}^{b}_{i},y^{b}_{i})$ (6)
where $\mathcal{L}_{ce}$ is cross-entropy loss [4], $\mathbf{p}^{b}_{i}$ donates the predicted probability of the $i$-th sample of the $b$-th task.
Inspired by  [38, 26], we also utilize contrastive learning [1] and data augmentation in the training of 3D FSCIL models for continual learning. Through the push and pull dynamics of contrastive learning, we optimize the extracted features, bringing them closer to the appropriate prototype while distancing them from erroneous ones. This strategy notably reduces ambiguity throughout the continual learning phase. Concurrently, data augmentation not only amplifies the efficacy of contrastive learning but also acts as a safeguard against overfitting in few-shot scenarios.
Specifically, we employ random rotations along the coordinate axes and random variations in the camera view distance as the augmentation function $f_{Aug}$. For the $i$-th sample of the $b$-th task, we train using the corresponding prototype $\mathbf{F}^{t}_{y^{b}_{i}}$ as a positive example and all other visible prototypes $\mathbf{F}^{t}_{res^{b}_{i}}$ as negative examples. We employ the InfoNCE loss [16] as contrastive learning loss. The contrastive learning loss $\mathcal{L}^{b}_{cont}$ for the $b$-th task can be calculated as follows:
$\mathcal{L}^{b}_{cont}=\frac{1}{|\mathcal{D}^{b}|}\sum_{i=1}^{|\mathcal{D}^{b}|}\sum_{j=1}^{N_{Aug}}\mathcal{L}_{InfoNCE}(\mathbf{f}^{g}_{i,j},\mathbf{F}^{t}_{y^{b}_{i}},\mathbf{F}^{t}_{res^{b}_{i}})$ (7)
where $N_{Aug}$ is the number of augmentations, and $\mathbf{f}^{g}_{i,j}$ is the global feature encoded after replacing $x^{b}_{i}$ with $f_{Aug}(x^{b}_{i})$.
The overall training loss $\mathcal{L}^{b}$ for $b$-th task can be calculated as follows:
$\mathcal{L}^{b}=\mathcal{L}^{b}_{cls}+\alpha\mathcal{L}^{b}_{cont}$ (8)
4 Benckmark for 3D FSCIL task
Studies on 3D FSCIL benchmarks are still in early stage. The sole benchmark introduced by Chowdhury et al. [3] has several notable limitations: 1) In the synthetic data to synthetic data (S2S) task set forth by Chowdhury et al. [3], the total number of classes is limited. This restricts the capacity of the model for incremental learning and makes it challenging to evaluate its effectiveness comprehensively. 2) For the synthetic data to real-scanned data (S2R) task, many analogous classes have been removed without clear justification. 3) Relying solely on accuracy as a metric means that incremental classes may not receive the emphasis they warrant. In light of these observations, we introduce a new benchmark: FSCIL3D-XL. Details of this benchmark will be discussed in the subsequent sections.
4.1 Task Setting
FSCIL3D-XL comprises two series of incremental tasks: S2S and S2R. The S2S task acts as a transitional and simulation-based task, primarily designed to assess the model’s capability to mitigate issues like overfitting or catastrophic forgetting. On the other hand, the S2R task is inherently more complex and has broader applicability. Given its larger domain gap and challenges with noise, it demands a more robust ability to generalize from limited samples.
S2S Task. Chowdhury et al. [3]’s dataset is constructed using a single dataset, encompassing only 55 classes. In our benchmark, we opt for ShapeNet [2] as our base dataset and ModelNet40 [36] as the incremental dataset. Our base task retains all the 55 classes from ShapeNet [2]. For the incremental tasks, we exclude 16 classes from ModelNet40 [36] that overlap with classes in the base task, and the remaining 24 unique classes from ModelNet40 are then evenly distributed across 6 incremental tasks. In contrast to Chowdhury et al. [3]’s work, our S2S task offers an increase in class count by over 40%, introducing a heightened level of challenge.
S2R Task. Chowdhury et al. [3]’s benchmark unjustifiably excludes certain classes that, while similar, are distinct in fact. For instance, semantically similar classes like “handbag” and “bag”, or shape-analogous classes like “ashcan” and “can”, are filtered out. By contrast, we retain all available classes in our benchmark. Recognizing the subtle differences between such similar classes is undoubtedly challenging, but we believe it is essential for a comprehensive evaluation. We continue to use ShapeNet as the base dataset and have selected CO3D [23] as the incremental dataset. From CO3D, we exclude 9 overlapping classes, resulting in 41 distinct classes designated for the incremental tasks.
For a more detail of the task settings in FSCIL3D-XL, please refer to Sec. 3 of Suppl..
4.2 Evaluation metrics
3D datasets present challenges distinct from their 2D counterparts. Firstly, there is a pronounced imbalance in the scale of different classes within 3D datasets. For instance, in ShapeNet, the “chair” class boasts over 1,000 training samples, whereas “birdhouse” has a mere 15. Secondly, incremental samples in 3D datasets, particularly those sourced from real-scanned data, tend to be more intricate. These novel classes, given their complexity, demand heightened attention in the FSCIL task. Relying solely on the accuracy metric, denoted as ${\rm Acc}$, would inadequately addresses these challenges. As a result, while we retain the 2D evaluation metrics ${\rm[Acc_{i}]_{B}}$ (accuracy of each session) and ${\Delta}$ [27] (relative accuracy dropping rate, where ${\rm\Delta=\frac{|Acc_{B}-Acc_{1}|}{Acc_{1}}}$), we also introduce new evaluation metrics specifically designed to address the aforementioned issues.
Macro accuracy. Macro accuracy (${\rm MAcc}$) indicates the generalization ability of the model, preventing overfitting in a small number of classes with a relatively large number of samples. It can be calculated using the following formula for $b$-th task:
${\rm MAcc_{b}}=\frac{1}{K_{b}}\sum_{i=1}^{K_{b}}Acc_{i}$ (9)
where $K_{b}=\sum_{i=1}^{b}|Y_{b}|$ donates the number of visible classes for the $b$-th task, and $Acc_{i}$ donates the accuracy of the $i$-th class.
Novel class accuracy. Novel class accuracy (${\rm NCAcc}$) indicates the ability of the model to learn new classes, preventing the model from over-focusing on base classes with a large number of samples. It can be calculated using the following formula for $b$-th task:
${\rm NCAcc_{b}}=\frac{|\{p_{i}^{b}|p_{i}^{b}=y_{i}^{b},p_{i}^{b}\in P_{b},y_{i}^{b}\in Y_{b}\}|}{|\mathcal{D}^{b}|}$ (10)
${\rm NCAcc}=\frac{1}{B}\sum_{b=1}^{B}{\rm NCAcc_{b}}$ (11)
where $P_{b}$ donates the predicted labels for test stage of $b$-th task and $B$ donates the number of task. F-Score. F-Score for FSCIL task (${\rm F_{FSCIL}}$, ${\rm F}$ for short): The network needs to be plastic to learn new knowledge from the current task, and it also needs to be stable to maintain knowledge learned from previous tasks. To ensure that the model is not overly biased towards either of these aspects, we refer to ${\rm F_{score}}$ and propose ${\rm F_{FSCIL}}$, which balances the plasticity and stability of the evaluation network. It can be calculated using the following formula:
${\rm F_{FSCIL}}=\frac{2\ {\rm Acc_{B}}\ {\rm NCAcc}}{{\rm Acc_{B}}+{\rm NCAcc}}$ (12)
where ${\rm Acc_{B}}$ denotes the accuracy of the ${\rm B}$-th (final) session.
Finally, the 3D dataset is still being refined, so a large number of new datasets will likely continue to be produced. We take this into account and design our benchmark to be more flexible and modular in adding datasets, i.e., the desired FSCIL dataset can be obtained by simply transmitting the parameters and datasets to our generator. Additionally, The benchmark will be open-resource.
5 Experiments
In this section, we first provide more detailed information about our models and experiments. Then, we present and analyze the results of the comparison experiments. Finally, we present our ablation experiment to verify the components of our model.
Table 1: Quantitative results on the S2S task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Pub. Year Acc. in each session $\uparrow$ Evaluation metrics
0 1 2 3 4 5 6 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
FACT [46] CVPR’22 82.6 77.0 72.4 69.8 68.4 67.7 67.3 41.7 18.5 51.5
48.0 44.7 42.0 39.9 38.2 37.1 36.5 34.0 23.9 35.2
BiDist [45] CVPR’23 89.6 87.7 86.2 84.7 83.8 83.6 82.3 35.0 8.1 49.1
82.5 80.3 77.3 75.2 72.5 71.2 69.3 38.0 16.0 49.1
Chowdhury et al.’s [3] ECCV’22 86.9 84.6 82.8 78.3 78.5 71.5 68.6 50.8 21.1 58.4
73.0 62.8 65.2 62.8 60.9 57.9 55.1 45.5 24.5 49.8
SimpleCIL-3D (ours) - 90.4 88.0 85.6 85.0 84.1 78.4 80.1 68.2 11.4 73.7
78.9 76.1 72.8 71.1 70.8 70.2 69.7 68.8 11.7 69.2
FILP-3D (ours) - 90.6 89.0 86.7 84.2 83.2 81.8 82.2 79.3 9.3 80.7
80.0 76.8 74.9 72.8 71.2 70.9 70.7 77.0 11.6 73.7
$Joint$ FILP-3D - 90.6 89.1 88.5 87.8 87.4 87.6 86.8 79.9 4.2 83.2
80.0 78.5 79.0 77.2 77.5 77.7 77.3 79.9 3.4 78.6
Table 2: Quantitative results on the S2R task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Acc. in each session $\uparrow$ Evaluation metrics
0 1 2 3 4 5 6 7 8 9 10 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
FACT [46] 82.4 77.2 74.5 73.1 71.3 70.4 67.2 65.2 63.8 61.8 59.9 59.8 26.2 27.4 36.4
48.6 41.4 39.7 36.8 35.5 33.6 31.2 29.5 28.4 27.2 25.8 25.9 30.8 46.7 28.1
BiDist [45] 89.4 54.0 54.7 56.4 57.0 55.9 56.3 52.9 52.3 51.7 50.8 50.1 47.2 43.9 48.6
81.8 52.8 50.1 46.2 48.3 46.1 44.7 41.8 41.8 39.8 40.0 39.7 41.9 51.5 40.8
Chowdhury et al.’s [3] 85.2 78.6 71.0 72.0 75.2 68.8 56.1 58.5 62.9 59.1 52.2 59.4 35.3 30.3 44.3
68.2 56.2 50.5 48.4 53.5 46.7 39.9 37.6 36.9 33.1 34.3 44.1 36.5 35.3 40.0
SimpleCIL-3D (ours) 89.2 86.7 83.5 81.7 79.4 79.6 78.6 70.4 72.1 71.7 70.1 71.2 49.6 20.2 58.5
78.9 77.3 75.9 73.7 70.1 66.5 64.4 60.9 59.4 58.5 54.5 56.3 49.0 28.6 52.4
FILP-3D (ours) 90.0 87.0 86.4 85.0 83.7 82.7 81.4 79.4 78.2 76.8 74.8 74.6 60.6 17.1 66.9
79.4 75.4 75.7 72.4 70.2 68.5 65.9 63.5 62.2 59.6 57.5 57.3 59.9 27.8 58.6
$Joint$ FILP-3D 90.0 89.2 89.0 88.4 88.1 87.7 87.2 86.9 85.9 84.6 83.1 83.1 54.9 7.7 66.1
79.4 77.2 76.3 75.9 75.5 73.5 73.3 73.4 70.2 69.3 68.2 68.2 60.0 14.1 63.8
5.1 Implementation Details
We choose CLIP’s ViT-B/32 [21] as our pre-trained model, replace its visual encoder with CLIP2Point’s pre-trained depth encoder [11] and adopt CLIP2Point’s proposed rendering approach. As for the 3D encoder, we use DGCNN [34] pre-trained on ShapeNet, which follows OcCo [32]. Dimensions of both image features and text features are 512. We use SVD [9] as our dimensionality reduction method. We extract principal components (242 out of 512) based on the base task, which retains 95% energy of the principal components. We set the temperature parameter of infoNCE $\tau$ and balance parameter $\alpha$ in Eq. 8 to 0.1 and 1.0 respectively. For training, we use ADAM weight decay optimizer [13]. We set the learning rate to $1\times 10^{-3}$ and the weight decay to $1\times 10^{-4}$. Training in the base task and incremental take 10 epochs and 20 epochs respectively. For incremental tasks, we randomly select 5 samples/classes for training and allow 1 exemplar/class from previous tasks to be used as memory. The batch size is 32.
5.2 Experimental Results
We conduct extensive experiments on FSCIL3D-XL’s S2S task (Table 1), FSCIL3D-XL’s S2R task (Table 2), and Chowdhury et al. [3]’s benchmark (Sec. 6 of Suppl.). To make a comprehensive comparison, we choose the following models: 1) $Joint$: Models are trained with samples from all currently visible classes, which can represent their upper bound in FSCIL tasks. 2) One 3D SOTA approach developed by Chowdhury et al. [3]. 3) Two 2D SOTA approaches FACT [46] and BiDist [45]. FACT uses manifold-mixup [31] and virtual prototypes to ensure forward compatibility. BiDist uses knowledge distillation to retain knowledge. We use official codes to reproduce these methods, keeping all the original parameters unchanged. The only difference is that we replace the CNN-based network with CLIP2Point’s depth encoder to allow 3D application. For visualization results, more comparison results (including ULIP [37] and PointCLIP [43]) and effect of our proposed metrics, please refer to Sec. 4,7,8 of Suppl.
Experiments on the S2S task of FSCIL3D-XL. Table 1 reports the results of the synthetic data to the synthetic data task in our benchmark. BiDist [45] additionally employs a weighted sum of distillation loss for optimization, along with a relatively higher loss weight assigned to it. Therefore, BiDist tends to preserve existing knowledge and sacrifice the learning capacity for new classes, leading to a relatively low $\rm NCAcc$. Moreover, the S2S task has a limited proportion of incremental data, BiDist can achieve top $\rm Acc$ in over half of the tasks by keeping old knowledge. Contrary to it, FILP-3D has a trade-off between the base classes and the novel classes. The results also show that our metrics are reasonable for the evaluation of 3D FSCIL.
Experiments on the S2R task of FSCIL3D-XL. Table 2 reports the results of the synthetic data to the real-scanned data task in our benchmark. Considering the large domain gap between synthetic data and real-scanned data, it is even harder to acquire knowledge from new classes based on previous knowledge of synthetic shapes. We notice that BiDist witnesses a catastrophic performance decline from task 1 to task 2, which further demonstrates that BiDist tends to retain old knowledge and sacrifice the acquisition of new knowledge. In contrast, SimpleCIL-3D has already outperformed the other three methods, which proves that prior knowledge from pre-training models is effective for few-shot incremental tasks. However, we notice that its $\rm NCAcc$ is much lower than ours, indicating that SimpleCIL-3D performs poorly in real-scanned data. With the Spatial Noise Compensator, our FILP-3D achieves a state-of-the-art performance, almost close to its upper bound.
Table 3: Ablation studies on the S2R task of FSCIL3D-XL. For no contrastive learning (CL) case, training loss $\mathcal{L}^{b}=\mathcal{L}^{b}_{cls}$
. RFE SNC CL 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$ 86.7 71.2 49.6 17.9 58.5 78.9 56.3 49.0 28.6 52.4 90.4 75.6 50.8 16.3 60.8 80.0 58.3 51.0 27.1 54.4 90.1 73.6 54.5 18.3 62.6 79.3 56.7 56.1 28.5 56.4 89.8 74.0 57.8 17.6 64.9 78.7 56.7 58.0 28.0 57.3 90.0 74.6 60.0 17.1 66.9 79.4 57.3 59.9 27.8 58.6
5.3 Ablation Studies
To evaluate the effectiveness of our proposed modules and components, we conduct several ablation experiments. We verify the RCS, the Redundant Feature Eliminator on the S2S task (Sec. 2 of Suppl.), and further verify the Redundant Feature Eliminator, the Spatial Noise Compensator and contrastive learning on the S2R task (Table 3).
Effect of Redundant Feature Eliminator. For no RFE case (line 1 and 3), we use cosine similarity between $\mathbf{f}^{g}$ and ${\rm\mathbf{F}}^{t}$ to calculate logits. One can notice that the model with RFE (line 2) can only slightly outperform SimpleCIL-3D (line 1) in $\rm NCAcc$. We analyze the reason as follows: The huge domain gap leads to some semantic features available in synthetic data can not be extracted efficiently on real-scanned data. Adding some real-scanned data samples in the principal component extraction stage as prior knowledge may effectively alleviate this symptom. Ablation study for RFE in the S2S task (Sec. 2 of Suppl.) can demonstrate a more reasonable comparison and prove the effectiveness of our method in solving redundant feature problems.
Effect of Spatial Noise Compensator. Let’s compare the results of the line 2 with the line 4. On the one hand, there is a slight decrease in $\rm Acc$, which indicates that the 3D module is not as capable as the multi-view model with 2D prior knowledge. It is suitable only for supplementing the information in the 3D FSCIL task. On the other hand, there is a significant increase in the incremental part of the evaluation metric, indicating that the 3D module can complement the global information that the multi-view model lacks and enhance the anti-interference ability.
6 Conclusion
In this paper, we proposed FILP-3D, a 3D few-shot class-incremental learning framework with pre-trained V-L models. Specifically, we introduced a V-L pre-trained model CLIP to the 3D FSCIL task. To guarantee that CLIP performs well in the 3D FSCIL task, we proposed a Redundant Feature Eliminator to eliminate redundant features without stretching semantic information and a Spatial Noise Compensator to complement noise information. To comprehensively evaluate the 3D FSCIL task, we further proposed a new benchmark FSCIL3D-XL, which retains all classes and introduces reasonable evaluation metrics. Extensive experiments demonstrate that our FILP-3D achieves state-of-the-art performance in all available benchmarks. Results from ablation studies further verify the effectiveness of the proposed components.
Supplementary Material
In the supplementary materials, we introduce a detailed discussion of RCS in Section 7 and conduct ablation experiments to demonstrate the effectiveness of RCS and RFE in Section 8 . We provide more details about the task settings of our proposed benchmark FSCIL-3D XL in Section 9 as a supplement to our proposed benchmark. We provide visualizations in Section 10 for a more intuitive comparison between our model and other models (or ablation results). We elaborate on the preprocessing details of the principal component $\mathbf{V}$ in Section 11 to demonstrate the details of our model. We present experiments conducted on Chowdhury et al.’s benchmark in Section 12 and comparison experiments with more methods in Section 13 to demonstrate the superiority of our model as well as the generality and effectiveness of our proposed modules. Finally, by analyzing the experimental results, we demonstrate how our proposed metrics offer a more comprehensive evaluation of 3D FSCIL models in Section 14.
7 Detail Discussion for RCS
$\mathbf{V}$ is the principal components extracted from the base task, $\mathbf{v}_{i}$ is the $i$-th normalized principal component. $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$ are the projection of $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$ onto the $\mathbf{V}$ vector space, formally expressed as the following equation:
$\mathbf{f}^{d}=\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}+\mathbf{R}^{d}$ (13)
$\mathbf{F}^{t}_{k}=\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{v}_{i}+\mathbf{R}^{t}_{k}$ (14)
where $M$ is the number of principal components, $\widetilde{\mathbf{f}}^{d}_{i}$ and $\mathbf{F}^{t}_{k,i}$ are the $i$-th dimension of $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}_{k}$, respectively, and $\mathbf{R}^{d}$ and $\mathbf{R}^{t}_{k}$ are the components of $\mathbf{f}^{d}$ and $\mathbf{F}^{t}_{k}$ which are not in the $\mathbf{V}$ vector space. $\mathbf{R}^{d}$ and $\mathbf{R}^{t}_{k}$ are low-variance components (due to the nature of principal component analysis) and thus are redundant components according to discussions in Section 3.4.1.
Each $\mathbf{v}_{i}$ is an orthogonal vector and orthogonal to vectors outside the $\mathbf{V}$ vector space, e.g., $\mathbf{R}^{d}$ and $\mathbf{R}^{t}_{k}$. Therefore $\forall i,j,k,\ i\neq j$ The following equation holds:
$\mathbf{v}_{i}\mathbf{v}^{T}_{j}=0,\ \mathbf{v}_{i}\mathbf{v}^{T}_{i}=1$ (15)
$\mathbf{R}^{d}\mathbf{v}_{i}^{T}=0,\ \mathbf{v}_{i}\mathbf{R}^{t^{T}}_{k}=0$ (16)
Then, we can derive the following equation:
$l_{k}$ $=\frac{\mathbf{f}^{d}\mathbf{F}^{t^{\mathrm{T}}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{(\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}+\mathbf{R}^{d})(\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{v}_{i}+\mathbf{R}^{t}_{k})^{T}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}\sum_{j=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,j}\mathbf{v}_{j}^{T}+\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$+\frac{\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\mathbf{v}_{i}\mathbf{R}^{t^{T}}_{k}+\sum_{i=1}^{M}\widetilde{\mathbf{F}}^{t}_{k,i}\mathbf{R}^{d}\mathbf{v}_{i}^{T}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\sum_{i=1}^{M}\widetilde{\mathbf{f}}^{d}_{i}\widetilde{\mathbf{F}}^{t}_{k,i}+\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
$=\frac{\widetilde{\mathbf{f}}^{d}\widetilde{\mathbf{F}}^{t^{T}}_{k}+\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}}{\|\mathbf{f}^{d}\|\|\mathbf{F}^{t}_{k}\|}$
Now we can observe that there are two terms in the numerator of this equation. $\widetilde{\mathbf{f}}^{d}\widetilde{\mathbf{F}}^{t^{T}}_{k}$ represents semantic information and should be retained, and $\mathbf{R}^{d}\mathbf{R}^{t^{T}}_{k}$ is the multiplication of redundant feature vectors that should be compressed to keep only semantic information. Thus, the logit of the class $k$ can be calculated in a modified way as follows to eliminate redundant features:
$\widetilde{l}_{k}=\frac{\widetilde{\mathbf{f}}^{g}\widetilde{\mathbf{F}}^{t^{T}}_{k}}{\|\mathbf{f}^{g}\|\|\mathbf{F}^{t}_{k}\|}.$ (17)
8 Extra Ablation Studies
Table 4: Ablation studies on the S2S task of FSCIL3D-XL
DR RCS 0 6 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
90.4 80.1 68.2 11.4 73.7
78.9 69.7 68.8 11.7 69.2
90.5 81.2 70.9 10.2 75.7
79.7 69.7 70.9 12.5 70.3
90.6 82.2 79.3 9.3 80.7
80.0 70.7 77.0 11.6 73.7
To further demonstrate the effectiveness of our designs, we provide more ablation results in Table 4. For no Dimensionality Reduction (DR) and no renormalized cosine similarity (RCS) case (line 1), we use cosine similarity between $\mathbf{f}^{d}$ and $\mathbf{F}^{t}$ to calculate logits. For with DR and no renormalized cosine similarity (RCS) case (line 2), we use cosine similarity between $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}$ to calculate logits. For with DR and renormalized cosine similarity (RCS) case (line 3), we use Eq.17 to calculate logits.
It can be observed that if DR is used without the RCS (line 2), The improvement in performance is limited. Conversely, the combination of DR and the RCS (line 3) leads to a significant improvement in performance, particularly in terms of $\rm NCAcc$. We can thus conclude that, without using the RCS, the process of dimensionality reduction leads to inappropriate stretching, which ultimately distorts the semantic information when eliminating redundant information. This phenomenon verifies the conclusion of our analysis that redundant information affects the classification.
We can observe a significant improvement in line 3 compared to line 1 (both in $\rm Acc$ and $\rm NCAcc$). This experimental result provides evidence that redundancy within PTMs can have a pronounced impact on classification accuracy and the effectiveness of RFE in mitigating this issue.
9 Details of FSCIL3D-XL
Dataset Task Name of Classes
ShapeNet 1 airplane, ashcan, bag, basket,
bathtub, bed, bench, birdhouse,
bookshelf, bottle, bowl, bus,
cabinet, camera, can, cap,
car, cellular telephone, chair, clock,
computer keyboard, dishwasher,
display, earphone, faucet, file,
guitar, helmet, jar, knife, lamp,
laptop, loudspeaker, mailbox,
microphone, microwave,
motorcycle, mug, piano, pillow,
pistol, pot, printer, remote control,
rifle, rocket, skateboard, sofa,
stove, table, telephone, tower,
train, vessel, washer
ModelNet 2 cone, cup, curtain, desk
3 door, dresser, flower pot, glass box
4 mantel, monitor, night stand, person
5 plant, radio, range hood, sink
6 stairs, stool, tent, toilet
7 tv stand, vase, wardrobe, xbox
Table 5: Details of S2S task setup.
Dataset Task Name of Classes
ShapeNet 1 airplane, ashcan, bag, basket,
bathtub, bed, bench, birdhouse,
bookshelf, bottle, bowl, bus,
cabinet, camera, can, cap,
car, cellular telephone, chair, clock,
computer keyboard, dishwasher,
display, earphone, faucet, file,
guitar, helmet, jar, knife, lamp,
laptop, loudspeaker, mailbox,
microphone, microwave,
motorcycle, mug, piano, pillow,
pistol, pot, printer, remote control,
rifle, rocket, skateboard, sofa,
stove, table, telephone, tower,
train, vessel, washer
CO3D 2 kite, keyboard, apple, plant
3 toaster, pizza, donut, parkingmeter
4 toybus, vase, baseballglove, couch
5 broccoli, hydrant, bicycle, toilet
6 toytrain, cup, banana, sandwich
7 book, mouse, hotdog, cellphone
8 baseballbat, umbrella,
toyplane, wineglass
9 tv, orange, toytruck, ball
10 stopsign, hairdryer,
backpack, remote
11 carrot, frisbee, cake, handbag
12 suitcase
Table 6: Details of S2R task setup.
9.1 S2S Task
We choose ShapeNet [2] as our base dataset and ModelNet40 [36] as the incremental dataset. Our base task retains all 55 classes from ShapeNet [2], with 42,001 training and 10,469 test samples. For the incremental tasks, we choose 24 non-overlapping (with the base task) classes, consisting of 1,339 test instances. Table 5 shows the detailed class divisions.
9.2 S2R Task
We use ShapeNet [2] as our base dataset (similar to the S2S task, with 42,001 training and 10,469 test samples) and then choose CO3D [36] as the incremental dataset. For the incremental tasks, we exclude 9 overlapping classes that overlap with the base task, resulting in 41 distinct classes designated for the incremental tasks, consisting of 2,928 test instances. Table 6 shows the detailed class divisions.
10 Visualization and Analysis
Figure 4: Visualization of experimental results.
The visualization results of some classes in the test set are presented in the first six figures of Figure 4. The first three figures present the visualizations for Chowdhury et al. [3]’s model (1), ours w/o RFE (2), and ours (3) in the S2S task. The next three figures similarly display the visualizations for Chowdhury et al. [3]’s model (4), ours w/o SNC (5), and ours (6) in the S2R task. The last two figures present the visualization results of all classes in the S2R test set, with ours w/o SNC on (7) and ours on (8). These visualizations are generated using t-SNE [30].
10.1 Comparison with Chowdhury et al. [3]’s
Comparing our model with Chowdhury et al. [3]’s, three observations can be concluded: 1) In our model, the intra-class samples appear more compact, indicating that samples within the same class are closer in the feature space. In contrast, the visualizations of Chowdhury et al. [3]’s model, using chair and bed as examples, exhibit a more scattered distribution. The lack of distinct clustering centers suggests its inability to extract more generalizable features. 2) In our model, there is a notable amount of blank space, indicating that other classes have sufficient feature space for representation. In contrast, the class distributions nearly fill the entire space in the visualizations of Chowdhury et al. [3]’s model, indicating a propensity for conflicts with unrepresented classes. 3) Considering the displayed classes in Figure 4, our model can accurately distinguish between bed and bus, chair and car, while Chowdhury et al. [3]’s model demonstrates confusion as indicated by the dashed boxes.
Based on the above observations, we can conclude that the features of our model surpass those extracted by Chowdhury et al. [3]’s model in aspects of both generalizability and discriminability.
10.2 Effect of RFE
Observing Figure 4 (2), we can notice that the bed and chair classes remain relatively dispersed compared to our model. This indicates that the model struggles to extract more generalizable features under the influence of redundant information, resulting in a more scattered distribution.
10.3 Effect of SNC
The distinction between ours w/o SNC (5) and ours (6) is not apparent in specific classes. Nevertheless, by comparing ours w/o SNC (7) with ours (8) in all classes, we can observe that noise greatly affects the discriminability of features (7). The inclusion of supplementary information through SNC effectively alleviates this issue (8).
11 Details of Pre-Processing
Figure 5: Overview of pre-processing
In the pre-processing stage, we render all training samples of the base task $[\mathbf{x}^{1}_{i}]_{M}$ to generate their depth maps $[{\rm\mathbf{D}}^{1}_{i}]_{M}$ and then embed them as depth features ${\rm\mathbf{F}^{2D}}\in\mathbb{R}^{MN\times C}$, where $M$ is the number of training sample in task 1, $N$ is the number of viewpoint and $C$ is the embedding dimension. We then use the depth features ${\rm\mathbf{F}^{2D}}$ to calculate principal components $\mathbf{V}\in\mathbb{R}^{P\times C}$ through SVD, as is shown in Figure 5, where $P$ is the number of principal component.
12 Experiments on Chowdhury et al. [3]’s benchmark.
Table 7: Quantitative results on Chowdhury et al. [3]’s ShapeNet2CO3D benchmark. Bold donates the best performance, $joint$ serves solely as an upper reference limit in Chowdhury et al. [3]’s work and is not involved in the comparison.
Method Acc. in each session $\uparrow$ ${\rm\Delta}\downarrow$
0 1 2 3 4 5 6 7 8 9 10
$Joint$ Chowdhury et al.’s [3] 81.0 79.5 78.3 75.2 75.1 74.8 72.3 71.3 70.0 68.8 67.3 16.9
FACT [46] 81.4 76.0 70.3 68.1 65.8 63.5 63.0 60.1 58.2 57.5 55.9 31.3
Chowdhury et al.’s  [3] 82.6 77.9 73.9 72.7 67.7 66.2 65.4 63.4 60.6 58.1 57.1 30.9
FILP-3D (ours) 85.4 81.3 80.1 77.9 75.1 73.8 72.3 70.6 69.3 67.1 65.6 23.2
Table 8: Extra quantitative results on the S2R task of FSCIL3D-XL. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Acc. in each session $\uparrow$ Evaluation metrics
0 1 2 3 4 5 6 7 8 9 10 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
ULIP [37] 86.3 85.6 81.7 74.0 71.7 68.1 67.6 64.5 59.5 58.4 55.2 57.5 56.4 33.4 56.9
85.4 82.0 78.2 70.4 66.9 63.3 61.9 58.1 54.8 52.1 49.2 50.3 57.0 41.1 53.4
PointCLIP [43] 87.5 83.2 79.5 74.9 75.6 74.2 71.1 71.7 67.9 66.5 64.3 66.1 54.1 24.5 59.5
74.2 72.1 68.8 62.3 60.1 57.9 54.5 55.1 50.9 50.8 47.5 48.7 53.2 34.4 50.9
PointCLIP++ 88.2 85.6 84.2 81.9 80.2 78.4 75.3 72.6 71.5 69.3 67.3 68.1 59.8 22.7 63.7
75.1 72.6 70.4 65.4 63.3 61.0 57.3 53.9 53.6 51.4 49.6 50.3 59.7 33.0 54.6
SimpleCIL-3D (ours) 89.2 86.7 83.5 81.7 79.4 79.6 78.6 70.4 72.1 71.7 70.1 71.2 49.6 20.2 58.5
78.9 77.3 75.9 73.7 70.1 66.5 64.4 60.9 59.4 58.5 54.5 56.3 49.0 28.6 52.4
FILP-3D (ours) 90.0 87.0 86.4 85.0 83.7 82.7 81.4 79.4 78.2 76.8 74.8 74.6 60.6 17.1 66.9
79.4 75.4 75.7 72.4 70.2 68.5 65.9 63.5 62.2 59.6 57.5 57.3 59.9 27.8 58.6
As shown in Table 7, FILP-3D outperforms the other two methods. After the final task, our accuracy only drops to 65.6%, which is even close to the upper bound of Chowdhury et al. [3]’s method. However, we note that the results in this benchmark cannot distinctly indicate whether the performance degradation is caused by the new class or the old classes. We cannot conclude more observations based on it.
Unfortunately, Chowdhury et al. [3]’s work does not provide any data partitions other than ShapeNet2CO3D. Consequently, we are unable to follow them and conduct further experiments and comparisons, such as incremental within a single dataset.
13 Extra Comparison Experiments
To explain why we choose 2D PTMs as the backbone rather than selecting a 3D backbone, we compare our approach with ULIP [37] (PointMLP [14] backbone). Besides, a considerable amount of research attempts to improve CLIP’s classification ability on point clouds, such as PointCLIP [43]. We compare our method with PointCLIP in the setting of 3D FSCIL and integrate our proposed modules onto PointCLIP to validate the effectiveness and generality of our proposed modules.
Comparing ULIP with our model, it’s evident that ULIP exhibits outstanding learning capabilities in novel classes. However, the large number of trainable parameters and severely inadequate data make it extremely challenging to avoid overfitting during the FSCIL process, ultimately resulting in the highest dropping rate $\rm\Delta$ and the worst performance, except the $\rm NCAcc$ metric. Consequently, we conclude that the combination of frozen 2D PTMs pre-trained on massive data along with a simple MLP adapter is more suitable for 3D FSCIL. Conversely, ULIP, aligned only with CLIP on a small amount of data, is not as suitable for the 3D FSCIL task.
Comparing PointCLIP with our model, we observe that the inter-view adapter proposed by PointCLIP exhibits similarly inadequate performance due to its complex structure. Moreover, to make a more fair comparison, we introduce PointCLIP++ (PointCLIP backbone + RFE + SNC; the contrastive learning cannot be similarly included due to PointCLIP’s inability to specify rendering distances). Experimental results demonstrate that PointCLIP++ outperforms PointCLIP significantly across various metrics, especially in $\rm NCAcc$. This further validates the generality and effectiveness of our proposed modules.
Table 9: Quantitative results on the S2S task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Pub. Year Acc. in each session $\uparrow$ Evaluation metrics
0 1 2 3 4 5 6 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
FACT [46] CVPR’22 82.6 77.0 72.4 69.8 68.4 67.7 67.3 41.7 18.5 51.5
48.0 44.7 42.0 39.9 38.2 37.1 36.5 34.0 23.9 35.2
BiDist [45] CVPR’23 89.6 87.7 86.2 84.7 83.8 83.6 82.3 35.0 8.1 49.1
82.5 80.3 77.3 75.2 72.5 71.2 69.3 38.0 16.0 49.1
Chowdhury et al.’s [3] ECCV’22 86.9 84.6 82.8 78.3 78.5 71.5 68.6 50.8 21.1 58.4
73.0 62.8 65.2 62.8 60.9 57.9 55.1 45.5 24.5 49.8
SimpleCIL-3D (ours) - 90.4 88.0 85.6 85.0 84.1 78.4 80.1 68.2 11.4 73.7
78.9 76.1 72.8 71.1 70.8 70.2 69.7 68.8 11.7 69.2
FILP-3D (ours) - 90.6 89.0 86.7 84.2 83.2 81.8 82.2 79.3 9.3 80.7
80.0 76.8 74.9 72.8 71.2 70.9 70.7 77.0 11.6 73.7
$Joint$ FILP-3D - 90.6 89.1 88.5 87.8 87.4 87.6 86.8 79.9 4.2 83.2
80.0 78.5 79.0 77.2 77.5 77.7 77.3 79.9 3.4 78.6
Table 10: Quantitative results on the S2R task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Acc. in each session $\uparrow$ Evaluation metrics
0 1 2 3 4 5 6 7 8 9 10 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
FACT [46] 82.4 77.2 74.5 73.1 71.3 70.4 67.2 65.2 63.8 61.8 59.9 59.8 26.2 27.4 36.4
48.6 41.4 39.7 36.8 35.5 33.6 31.2 29.5 28.4 27.2 25.8 25.9 30.8 46.7 28.1
BiDist [45] 89.4 54.0 54.7 56.4 57.0 55.9 56.3 52.9 52.3 51.7 50.8 50.1 47.2 43.9 48.6
81.8 52.8 50.1 46.2 48.3 46.1 44.7 41.8 41.8 39.8 40.0 39.7 41.9 51.5 40.8
Chowdhury et al.’s [3] 85.2 78.6 71.0 72.0 75.2 68.8 56.1 58.5 62.9 59.1 52.2 59.4 35.3 30.3 44.3
68.2 56.2 50.5 48.4 53.5 46.7 39.9 37.6 36.9 33.1 34.3 44.1 36.5 35.3 40.0
SimpleCIL-3D (ours) 89.2 86.7 83.5 81.7 79.4 79.6 78.6 70.4 72.1 71.7 70.1 71.2 49.6 20.2 58.5
78.9 77.3 75.9 73.7 70.1 66.5 64.4 60.9 59.4 58.5 54.5 56.3 49.0 28.6 52.4
FILP-3D (ours) 90.0 87.0 86.4 85.0 83.7 82.7 81.4 79.4 78.2 76.8 74.8 74.6 60.6 17.1 66.9
79.4 75.4 75.7 72.4 70.2 68.5 65.9 63.5 62.2 59.6 57.5 57.3 59.9 27.8 58.6
$Joint$ FILP-3D 90.0 89.2 89.0 88.4 88.1 87.7 87.2 86.9 85.9 84.6 83.1 83.1 54.9 7.7 66.1
79.4 77.2 76.3 75.9 75.5 73.5 73.3 73.4 70.2 69.3 68.2 68.2 60.0 14.1 63.8
14 Effect of our proposed metrics
14.1 S2S task
By observing the experiment results presented in Table 9, we notice a substantial decline in Chowdhury et al.’s [3] performance in terms of $\rm MAcc$ when learning incrementally, coupled with poor performance in the $\rm NCAcc$ metric. This signifies their inability to effectively acquire new knowledge. Solely relying on micro-Acc. and dropping rate $\Delta$, we are unable to identify the aforementioned issues. Similarly, without $\rm NCAcc$, Bidist’s inability to effectively learn new classes also can not be revealed.
14.2 S2R task
By observing the experiment results presented in Table 10, we notice a catastrophic decline in FACT [46]’s performance in terms of $\rm MAcc$ when learning real-scanned data. This may be because The feature space trained by the manifold-mixup [31] method is relatively fragile in the aspect of structure, thus it is easy to overfit the noise information during the incremental process, especially for real-scanned data. Solely relying on metrics proposed by Chowdhury et al., we are unable to identify the aforementioned issues and then perform the analysis. Similarly, without $\rm NCAcc$, Chowdhury et al.’s terrible performance in learning real scanned new classes also can not be noticed.
Consequently, our proposed metrics, $\rm MAcc$ and $\rm NCAcc$, hold significant importance in comprehensively evaluating and analyzing the 3D FSCIL model. Moreover, ${\rm F_{FSCIL}}$, serving as a metric that balances learning new classes and not forgetting old ones, can objectively evaluate the performance of a continual learning model as much as possible.